If the debt of a loan increases by 37% in 6 years, what is the loan's equivalent annual rate?
Answer #1

Ans : Future value = 137 %

Present value = 100%

number of years(n) = 6 years

rate of interest (r) = ?

Future value = Present value * (1+r)n

137 = 100 * (1+r)6

1.37 = (1+r)6

(1+r) = (1.37)1/6

1+r = 1.053869

r = 5.3869 %